The correct size of a nickel is 21.21 millimeters. Based on that, the data can be summarized into the following table: Too Small Too Large Total Low Income 23 17 40 High Income 24 11 35 Total 47 28 75 Based on this data: (give your answers to parts a-c as fractions, or decimals to at least 3 decimal places. Give your to part d as a whole number.) a) The proportion of all children that drew the nickel too small is: Assume that this proportion is true for ALL children (e.g. that this proportion applies to any group of children), and that the remainder of the questions in this section apply to selections from the population of ALL children. b) If 7 children are chosen, the probability that exactly 4 would draw the nickel too small is: c) If 7 children are chosen at random, the probability that at least one would draw the nickel too small is: d) If 80 children are chosen at random, it would be unusual if more than drew the nickel too small
Answer #1

a) The proportion of all children that drew the nickel too small is =47/75 =0.627

b)

probability that exactly 4 would draw the nickel too small is =7C4(47/75)4(28/75)3 =0.281

c)

probability that at least one would draw the nickel too small is =1-(1-47/75)7 =0.999

d)

(please revert for this part as it is not given how much in more than section)